OpenStudy (anonymous):

Medal and fan Solve for x: 5x + 15 = 28 forty-three over five thirteen over five 13 43

OpenStudy (anonymous):

What do you think your first step is in solving this equation?

OpenStudy (anonymous):

subtracting

OpenStudy (anonymous):

Yeppers. Subtracting 15 from each side. What equation do you get if you do that?

OpenStudy (anonymous):

what after subtracting from both sides I got 13 and 0

OpenStudy (anonymous):

what about the 5x

OpenStudy (anonymous):

Ah, common mistake. Here's what you're actually doing:

OpenStudy (anonymous):

Most times, people think you're just taking 15 and subtracting 15, which would give you 0. This isn't the case. You're actually subtracting 15 from the entire SIDE. Here's what it actually looks like.\[5x+15=28\]\[5x+15-15=28-15\] It's true that the fifteens will kind of "cancel each other out," but you're still left with a 15x on the left side, not just zero. \[5x+15=28\]\[5x+15-15=28-15\]\[5x=13\]

OpenStudy (anonymous):

*That should have said 5x, not 15x

OpenStudy (anonymous):

Let me know if that doesn't make sense and I can clarify it some more

OpenStudy (anonymous):

what do I do with the 5x

OpenStudy (anonymous):

if I'm solving for x

OpenStudy (texaschic101):

divide both sides by 5

OpenStudy (texaschic101):

to get rid of a number on one side to put it on the other side, or to cancel it out...do the opposite. If it shows add, you subtract and vice versa. If it shows multiply, you divide...and vice versa

OpenStudy (anonymous):

So you want x, not 5x. Eventually you want the equation to look like x = # The best way to do this is to get rid of the 5. Now, when we write 5x, we actually mean 5 times x. To get rid of multiplication by 5, we divide by 5. So divide 5x by 5. But what we do to one side, we must do to the other, so we divide the other side by 5. \[5x=13\]\[\frac{5x}{5}=\frac{13}{5}\]\[x=\frac{13}{5}\]

OpenStudy (texaschic101):

but don't forget, what you do on one side of the equal sign, you have to do on the other side

OpenStudy (texaschic101):

good explaining jabberwock :)

OpenStudy (anonymous):

lol, sorry, I was mid-type when you started replying.

OpenStudy (anonymous):

I figured I'd just finish. Hopefully either your way or mine will make sense :) (although it looks like we're approaching it the same way)

OpenStudy (texaschic101):

great minds think alike..lol

OpenStudy (anonymous):

so the answer is 13 divided by 5

OpenStudy (anonymous):

well the answer of 13 divided by 5

OpenStudy (texaschic101):

13/5......or if you need it in mixed number form, it is 2 3/5 (2 and 3/5)

OpenStudy (anonymous):

Yep, either way would work. Looks like you have a multiple-choice question though.

OpenStudy (anonymous):

ok thank both of you who do I give the medal and fan to

OpenStudy (anonymous):

I'm not here much. Give it to texas

OpenStudy (texaschic101):

jabberwock

OpenStudy (texaschic101):

give it to jabberwock

OpenStudy (anonymous):

How about I give a fan to @jabberwock and a medal to @texaschic101 seems fair @jabberwock has 3 medals and @texaschic101 has one

OpenStudy (texaschic101):

jabberwock helped you out more on this one. Besides, I do not do this for medals. If jabberwock keeps answering questions like this, she/he will be at 99 before you know it.

OpenStudy (anonymous):

you can still give him a medal @texaschic101

OpenStudy (anonymous):

bi everyone